OpenStudy (anonymous):

1/R+1/3R=1/Rtot solve for Rtot

OpenStudy (anonymous):

\[\frac{1}{R}+\frac{1}{3R}=\frac{1}{R_{tot}}\] this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

start by simplifying \[\frac{1}{R}+\frac{1}{3R}=?\] then multiply the answer by 1 so that the numerator will be equal to 1 \[1=\frac{2}{2}=\frac{\frac{1}{2}}{\frac{1}{2}}\] then the denominator will be R_tot

OpenStudy (anonymous):

can you solve I do not understand

OpenStudy (anonymous):

ok then ill just simplify the first part \[\frac{1}{R}+\frac{1}{3R}\] \[1*\frac{1}{R}+\frac{1}{3R}\] \[\frac{3}{3}*\frac{1}{R}+\frac{1}{3R}\] \[\frac{3}{3R}+\frac{1}{3R}\] \[\frac{4}{3R}\]

OpenStudy (anonymous):

good luck with the rest

OpenStudy (anonymous):

4/3R=1/Rtot 4Rtot=3R?

OpenStudy (anonymous):

yes keep going

OpenStudy (anonymous):

emmm I dont know how to keep going

OpenStudy (anonymous):

there is 2 diffrent variables

OpenStudy (anonymous):

thats fine, just solve for R_tot in terms of R

OpenStudy (anonymous):

so Rtot= 3R/4?

OpenStudy (anonymous):

yea thats it

OpenStudy (anonymous):

Thanks so much !!!!!!

OpenStudy (anonymous):

no prob